# Deanonymizing Ethereum Validators: The P2P Network Has a Privacy Issue

### v1

@@ Appendix A Ethical Considerations
Respect for Persons. The principle of respect for persons requires that individuals’ autonomy and privacy are protected, and informed consent is obtained whenever necessary. Validators are logical entities identified through pseudonyms, which by design should make it so that they cannot be linked to humans. Nonetheless, we ensured that no identifiable information about any entities running Ethereum validators was exposed. We minimized the use of identifiable data, by using it purely to verify the validity of our results. Moreover, the analysis was done using publicly available data from the Ethereum peer-to-peer protocols.
Beneficience. We systematically assess risks by identifying key stakeholders, in order to maximize probable benefits while minimizing probable harm. First, staking entities that run Ethereum validators might suffer monetary harm, if an attacker were to exploit the vulnerability we expose. In order to minimize the likelihood of this scenario, we publish neither our code nor the deanonymized validator dataset. We further reached out to some larger entities, informing them of our findings. We were also careful to only run standard Ethereum clients to collect data. In particular, we avoided rewriting client code to make deanonymization more efficient, as this might have caused additional load on current Ethereum nodes. Instead, our clients most likely had a net positive impact, by contributing in forwarding P2P messages for the duration of our experiment.
@@
Appendix B Beacon Network Crawler
Below we detail our measurements of the Ethereum Beacon network via our own crawler implementation. A node in the Beacon network keeps a peer table of nodes they have heard about, which they reference when they need to make more peer connections. The peer table is based on the Kamdemlia hash table [50] where each node stores Ethereum Node Records (ENRs) into buckets of size 16 with each bucket containing nodes whose ids (a 256-bit unique identifier computed from the ENR) has an XOR distance of $i$ from itself for $2560\leq i<$ 256. Thus each bucket contains nodes whose ids share the first $i$ bits with itself. In practice, this means only around the first 16 buckets contain records.
(a) FR
(b) SO
(c) VA
(d) ZH
Figure 10: Hisogram showing the total connection time per peer for each of the four Rainbow nodes. The vertical red dashed line indicates the cutoff for long connections, i.e., we have enough data to attempt deanonymization of peers right of the line. Notice the logarithmic scale on the y-axis.
Nodes periodically update their peer tables using the discv5 peer discovery protocol [27]. In the discovery protocol, a node sends a FINDNODE request to some peer, with an input that indicates the peer table entry (the Kademlia “bucket") where the peer should send nodes from (i.e., the distance corresponding to the bucket). The peer responds with a NODES message containing up to 16 ENRS of nodes in that bucket or nearby buckets if the bucket does not contain enough entries (more common for smaller distances). We can use these messages to enumerate a node’s peer table by sending FINDNODE requests for each bucket of the node. We do this for all nodes we hear about in the network.
Figure 9 shows the count of nodes on the Beacon network discovered in our crawler each day that are reachable (i.e., respond to an in-protocol network ping) each hour. We note that the in-protocol ping messages use the node’s Ethereum Node Record (ENR) value, and some nodes do not respond if the ENR used is not up-to-date. Since the ENR changes if any of the values encoded change (e.g., a change in subnet subscriptions), ENRs change with some frequency. Therefore, the gradual drop of responses over time we see in Figure 9 can be explained by both some general network churn and ENRs learned in the previous crawl becoming outdated.
Appendix C Peer Connection Times
In Figure 10, we show a histogram of the total connection length our four Rainbow nodes maintained with their peers over the three days of our data collection. Unexpectedly there are many peers with whom our Rainbow nodes only maintain short connections and considerably fewer peers with whom they maintain long connections. Running the nodes for longer likely leads to more long connections.
Appendix D Staking Pool Validator Distribution
In Figure 11 examines the distribution of validators in large liquid staking services. The histogram shows the number of validators hosted by peers from the top five staking pools (four of which are deanonymized). While some peers host fewer than 100 validators, most peers in these pools host over 100, highlighting the economic benefits of staking pools.

### v2

@@ 12 Ethics Considerations
Respect for Persons. The principle of respect for persons requires that individuals’ autonomy and privacy are protected, and informed consent is obtained whenever necessary. Validators are logical entities identified through pseudonyms, which by design should make it so that they cannot be linked to humans. Nonetheless, we ensured that no identifiable information about any entities running Ethereum validators was exposed. We minimized the use of identifiable data, by using it purely to verify the validity of our results. Moreover, the analysis was done using publicly available data from the Ethereum P2P protocols.
Beneficience. We systematically assess risks by identifying key stakeholders, in order to maximize probable benefits while minimizing probable harm. First, staking entities that run Ethereum validators might suffer monetary harm, if an attacker were to exploit the vulnerability we expose. In order to minimize the likelihood of this scenario, we publish neither our code nor the deanonymized validator dataset (see Section 13). We further reached out to some larger entities, informing them of our findings. We were also careful to only run standard Ethereum clients to collect data. In particular, we avoided rewriting client code to make deanonymization more efficient, as this might have caused additional load on current Ethereum nodes. Instead, our clients most likely had a net positive impact, by contributing in forwarding P2P messages for the duration of our experiment.
@@
13 Open Science
As discussed in Section 12, we believe that our collected data is sensitive and should not be shared. Therefore, we only present aggregated data and refrain from any analysis targeting single entities. We also suggest keeping our Rainbow client private, until appropriate countermeasures have been implemented. Therefore, we refrain from making them publicly available.
Acknowledgments
We thank the anonymous USENIX Security 2025 reviewers for their useful suggestions. L. Kiffer contributed to this project while partially supported by the armasuisse Science and Technology CYD Distinguished Postdoctoral Fellowship. We also thank Amazon Web Services for a gift of AWS credit used towards the collection of our data.
Appendix A Heuristics
To test the robustness of our heuristics, in Tables 7 and 8 we investigate the impact of each condition on the effectiveness and accuracy of our deanonymization technique.
Table 7 analyzes how many validators we locate in total, as well as excluding those associated with service providers in columns one and two. We further use the verification method introduced in Section 7.2 to indicate how many non-unique validators we find and what proportion this makes up of the validators (excl. service providers). Each of the conditions is tested separately in the respective subtables.
In Table 8, we extend the robustness testing of our parameter choices by employing the verification metrics introduced in Section 7.1. We measure the proportion of our peers who are assigned a validator set that is consistent, inconsistent, or unknown (validators whose labels do not pass consistency checks). We also present the same statistics for validators as opposed to peers, e.g., how many validators are part of a consistent validator set. Again, each of the conditions is tested separately in the respective subtables.
C1 C2 C3 C4 validators validators (excl. service providers) non-unique validators % non-unique validators
\ 1 10 1 355,493 271,260 35,676 13.15
0.30 1 10 1 262,408 165,083 16,487 9.99
0.90 1 10 1 252,895 161,057 16,172 10.04
1.00 1 10 1 208,481 125,578 7,150 5.69
(a) C1
C1 C2 C3 C4 validators validators (excl. service providers) non-unique validators % non-unique validators
0.90 \ 10 1 270,242 181,494 20,117 11.08
0.90 1 10 1 252,895 161,057 16,172 10.04
(b) C2
C1 C2 C3 C4 validators validators (excl. service providers) non-unique validators % non-unique validators
0.90 1 \ 1 893,509 885,864 473,193 53.42
0.90 1 20 1 376,088 270,441 29,111 10.76
0.90 1 10 1 252,895 161,057 16,172 10.04
0.90 1 5 1 150,219 110,896 11,906 10.74
(c) C3
C1 C2 C3 C4 validators validators (excl. service providers) non-unique validators % non-unique validators
0.90 1 10 \ 256,725 162,168 16,350 10.08
0.90 1 10 1 252,895 161,057 16,172 10.04
0.90 1 10 4 159,494 136,350 15,308 11.23
(d) C4
Table 7: Number of validators located by each Rainbow node and overall. The first column indicates the total number of validators, the second column excludes validators from peers we identify to be P2P service providers (see Section 7.1), and the third column indicates the validators with a non-unique mapping to an IP port combination (see Section 7.2). We repeat the analysis for various parameter configurations of our heuristics, testing the parameter for each condition individually. Our chosen configuration is highlighted in bold.
A.1 Condition C1
Condition C1 ensures that the proportion of non-backbone attestations remains sufficiently large. The parameter $c1$ dictates how closely the observed ratio must align with the expected ratio. For example, when $c1$ is set to 0.9, the observed proportion must reach at least 90% of the expected value to satisfy the condition.
Our analysis compares scenarios without condition C1 to those with three different $c1$ values (0.3, 0.9, and 1.0). We observe that omitting condition C1 results in a nearly one-third increase in the total number of validators and an almost 50% increase when excluding service providers (see Table 7(a)). Alarmingly, the number of non-unique validators more than doubles, leading to a higher overall proportion of non-unique validators. Turning to Table 8(a), we find that without condition C1, the consistency of validator sets diminishes significantly, further highlighting the importance of this condition as part of our deanonymization methodology.
For the remaining values we test for C1, we observe smaller differences. Even with $c1$ set to 0.3, we observe similar numbers across the board as with $c1$ set to 0.9 (the parameter we choose). In general, we find slightly fewer validators with $c1$ set to 0.9, but observe that the validator sets we locate are more consistent. For the even more restrictive configuration ($c1=1.0$), we see that we lose a significant number of validators and have a smaller proportion of non-unique validators (see Table 7(a)). At the same time, the validator sets do not grow more consistent (see Table 8(a)).
In conclusion, our analysis demonstrates that condition C1 is essential for reducing likely false positives, as evidenced by the significant increase in non-unique validators and inconsistency when C1 is absent. We select $c1=0.9$ as it strikes a balance between being restrictive and accommodating minor deviations from the expected ratio, allowing for small discrepancies without compromising the overall consistency.
A.2 Condition C2
Condition C2 ensures that the node is not subscribed to all subnets. If a node were subscribed to all subnets, it would be impossible to receive non-backbone attestations from that node, rendering condition C1 ineffective.
We perform the analysis both without and with (indicated by a 1 in Table 7(a) and Table 8(b)) condition C2. Without C2, we observe a slight rise in the number of validators, but at a small cost in the proportion of non-unique validators (see Table 7(b)) and a slight increase in the number of inconsistent validator sets on peers (see Table 8(b)). Thus, we use condition C2 to avoid having to rely on less obvious signals (i.e., number of messages as opposed to type of message.
C1 C2 C3 C4 peers validators
consistent inconsistent unknown consistent inconsistent unknown
\ 1 10 1 74.32% 19.92% 5.75% 47.15% 55.73% 1.50%
0.30 1 10 1 92.56% 1.65% 5.79% 61.84% 41.50% 2.10%
0.90 1 10 1 93.75% 0.92% 5.33% 63.67% 39.67% 2.24%
1.00 1 10 1 93.38% 0.77% 5.85% 58.18% 44.24% 2.93%
(a) C1
C1 C2 C3 C4 peers validators
consistent inconsistent unknown consistent inconsistent unknown
0.90 \ 10 1 93.73% 0.97% 5.30% 65.21% 37.98% 2.11%
0.90 1 10 1 93.75% 0.92% 5.33% 63.67% 39.67% 2.24%
(b) C2
C1 C2 C3 C4 peers validators
consistent inconsistent unknown consistent inconsistent unknown
0.90 1 \ 1 66.17% 27.70% 6.13% 7.70% 94.71% 0.23%
0.90 1 20 1 88.37% 6.29% 5.33% 34.88% 66.86% 1.01%
0.90 1 10 1 93.75% 0.92% 5.33% 63.67% 39.67% 2.24%
0.90 1 5 1 94.03% 0.65% 5.32% 78.19% 23.25% 3.20%
(c) C3
C1 C2 C3 C4 peers validators
consistent inconsistent unknown consistent inconsistent unknown
0.90 1 10 \ 93.76% 0.92% 5.33% 63.21% 40.13% 2.24%
0.90 1 10 1 93.75% 0.92% 5.33% 63.67% 39.67% 2.24%
0.90 1 10 4 93.83% 0.83% 5.34% 86.28% 10.74% 3.65%
(d) C4
Table 8: Percentage of peers with a consistent validator set and validators belonging to a consistent validator set across our four Rainbow nodes as well as overall. Consistency is defined in Section 7.1. We repeat the analysis for various parameter configurations of our heuristics, testing the parameter for each condition individually. Our chosen configuration is highlighted in bold.
A.3 Condition C3
With condition C3 we ensure that we get at least every $c3$-th message from a peer for a specific validator that we locate on the peer. This is important to ensure that we receive enough data and maintain a good connection with the peer.
We test the condition C3 by repeating the analysis without C3 and by using three different parameter configurations (i.e., 20, 10, and 5). The analysis reveals that in the absence of C3 the number of validators is more than three-fold and the number of non-unique validators skyrockets (see Table 7(c)). Further, the consistency of validator sets on peers decreases significantly (see Table 8(c)). This underscores the importance of having a good enough connection as a prerequisite for our deanonymization methodology.
For the various parameters for $c3$ we test, the differences are less extreme but still significant. With $c3=20$, the least restrictive parameter combination we test, we increase the validator count by almost one-third without significant increases in terms of the proportion of non-unique validators, but with a noticeable decrease in the proportion of consistent peers in comparison to our chosen parameter combination ($c3=10$). When comparing $c3=10$ to $c3=5$, we observe that with $c3=10$ we find significantly more validators but experience almost no losses in terms of the proportion of non-unique validators or the proportion of peers with consistent validator sets. We only observe that the proportion of validators in consistent validator sets decreases. However, this is due to more peers being associated with P2P service providers than otherwise (see Section 7.1). As these are excluded in our analysis, there is no significant negative impact of choosing the slightly less restrictive parameter $c3=10$.
To summarize, condition C3 ensures that we only try to deanonymize validators on a peer if we were able to maintain a good connection. Otherwise, false positives might become more likely, as indicated by the rise in non-unique validators and inconsistent validator sets in the absence of condition C3.
A.4 Condition C4
Our final condition C4 ensures that we receive significantly more messages from a validator we map to a peer than for the remaining validators for which we receive messages from the same peer. In particular, we require that the number of messages received for a validator we deanonymize on a peer is $c4$ standard deviations more than for the mean across all validators for which we receive messages from the peer. This heuristic is an extra precautious measure to ensure that there is a separation in terms of the number of messages between validators mapped to the peer and those not.
To test this condition, we perform the analysis without C4 and for two values for $c4$, namely 1 and 4. We find that the results appear largely similar for all configurations in terms of the proportion of non-unique validators (see Table 7(d)) and consistency (see Table 8(d)). The main difference is that for the most restrictive $c4$ value of 4, we locate significantly fewer validators. Thus, we choose $c4=1$ for our parameter configuration.
Note that even though this heuristic has the smallest impact, as revealed by our testing, we have the heuristic in place as a measure of caution.
A.5 Fanout Membership and False-Negatives
To deanonymize a peer, we rely on receiving attestations that originate from that peer (i.e., attestations signed by a validator running on that peer’s machine) on subnets where the peer is currently not participating as a backbone. To receive these non-backbone attestations, the peer must have added us to their fanout for that subnet. In our deanonymization approach, we consider only peers for which we are included in several fanouts (controlled by Condition C3). We note that if we ever participate in any fanout – i.e., by receiving at least one non-backbone attestation [^21] This could also happen during dynamic subset membership, which implies a validator that is aggregating.– this peer is considered as potentially hosting a validator, and is thus either deanonymized, or included in the "rest" column of Table 3 (those peers that would need more careful examination). We, therefore, only consider peers who have never sent a non-backbone attestation (and are not in all 64 subnets) as running no validators.
We now argue the unlikelihood of never receiving a non-backbone attestation from a peer running a validator based on the gossip protocol specs [53]. [^22] We use the default parameters of the Prysm client but note other implementations have similar parameters. When we first connect to the peer, they add us to all fanouts that are not at capacity (default fanout size is 8) that we are a backbone in (i.e., all of them). Thus, a peer participating in 2 subnets will have 62 fanouts. Assuming they are running with a default peer cap of 50-100 peers all on 2 subnets, in expectation each fanout would have less than 4 peers (be half empty), thus we would be added to all of them. Even allowing for some of their peers running on all 64 subnets, [^23] Note that less than 1% of our peers run on all subnets (see Table 3). this would still leave most fanouts not at capacity in expectation. Thus, when we connect for the first time to a peer, we are quite likely added to at least one fanout. A peer running with non-default parameters may run with higher peer counts and thus have full fanouts. These could lead to false negatives in our approach. We note that the longer we maintain a connection to a peer, the stronger the likelihood of eventually being added to one fanout long enough to receive at least one non-backbone attestation.